the faces of a cube are randomly and independently painted either red or blue with equal likelihood. What is the probability that the cube has all blue faces ? Express your answer as a common fraction ?
I tried 1- P(0)= 1 - (6c0)/64 is this correct ?

Answer by stanbon(75887)

the faces of a cube are randomly and independently painted either red or blue with equal likelihood. What is the probability that the cube has all blue faces ? Express your answer as a common fraction ?
----
Binomial Problem::
n = # of faces = 6
p(red) = 1/2 = p(blue)
-----------------
P(x = 6 blue) = (1/2)^6 = 1/64
-------------------
Cheers,
Stan H.
-------------